OpenStudy (josedavid):

Write an equation of a line that passes through the point (2, 3) and is parallel to the line y = 3 over 2x + 5. y = negative 2 over 3x y = negative 2 over 3x − 7 y = 3 over 2x y = 3 over 2x + 7

OpenStudy (photon336):

@josedavid a parallel lines have the same slope

OpenStudy (josedavid):

i think yes

OpenStudy (photon336):

So we can re-write this as \[y = \frac{ 3 }{ 2x+5}\] \[y_{2} = \frac{ 3 }{2x }\] now it must pass through point (2,3) How do you think we can solve this part?